A Bibliography of Supercomputing ’97

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

05 September 2013 Version 1.07

Abstract

This bibliography records articles presented at the Supercomputing ’97 conference.

Title word cross-reference

N [BS97b].

-Body [BS97b].

Access [MR97]. ACM/IEEE [ACM97]. Active [LMC97]. Adaptive [BS97b, PBEK97, SdG^{+}97]. Aeroacoustic [NOL97]. Air [LSA97]. Algebra [ABE^{+}97]. Algorithm [Fie97]. Algorithms [PBEK97]. Analysis [ABGS97, TNM^{+}97]. Application [KKCB97, KG97]. Application-level [KG97]. Applications [KPS97, LLWB97]. Architecture [GBF^{+}97, KG97, Lei97, SB97]. Architectures [BW97]. ASCI [LLWB97]. Assimilation [LEG^{+}97]. Atmosphere [TFS^{+}97]. Atmosphere-ocean [TFS^{+}97].

Balancing [HK97]. based [EGCL97, KK97, KG97, NOL97]. Benard [CHMS97]. Benchmarks [HLK97]. block [HJG^{+}97]. Blocks [SdG^{+}97]. Body [BS97b].

Cache [HLK97]. Cache-coherent [HLK97]. Calculations [LLC^{+}97]. California [ACM97]. Carbon [SB97]. Cases [TOK97]. centric [LLWB97]. Characteristics [Jet97]. Characterization [LLWB97]. Checkpointing [CLP97, NMM97]. Circulation [WKC97]. Classification [MR97]. client [TNM^{+}97]. Climate [TFS^{+}97]. CLIP [CLP97]. Cluster [LMC97, TOK97]. Clusters [BS97a]. CM [NOL97]. Code [KPS97]. coherent [HLK97]. Collaborative [Lei97]. Collector [ETY97]. Common [PBEK97]. Communication [BS97a, CCvK97, CC97]. Compilers [LZ97]. Compiling [KPS97, RMCKB97]. Computation [HJ97, KG97]. Computers [EGCL97]. Computing [ACM97, GBF^{+}97, LEG^{+}97, MT97, TNM^{+}97]. Concept [HJG^{+}97]. Conference [ACM97]. Conquer [de 97]. Cooperation [ABC^{+}97]. CORBA [KG97]. CORBA-based [KG97]. CRAY [LLC^{+}97].

Data [Joh97, LEG^{+}97, NSC97, PBEK97, SdG^{+}97, TOK97]. Database [TOK97]. Decision [LSA97, TOK97]. Decks [NOL97]. Dense [Fau97]. Design [ABE^{+}97, Lei97, LSA97]. Diffusion [Fau97]. Distributed [HJ97, Joh97, KG97, Lei97, NMM97, BW97]. Divide [de 97]. duration [TFS^{+}97]. During [VG97]. Dusty [NOL97]. Dynamic [HK97, WM97]. Dynamics [SB97].

Effects [BS97a]. Efficient [BS97b]. End [BS97a, GBF^{+}97]. Environment [MFGD97]. Environments [Jet97, Lei97]. Equation [LZ97]. Evaluating [CCvK97]. Evaluation [KKCB97, LZ97, LLWB97]. Experiment [KM97]. Exploiting [MR97].

Factorization [KK97]. fetching [VG97]. fidelity [HJG^{+}97]. File [BDK97]. Flexible [Lei97]. Flow [CHMS97]. FM [CC97]. FM-QoS [CC97]. FOAM [TFS^{+}97]. Fock [KKCB97]. Forecasting [WSP97]. Fortran [NOL97, RMCKB97]. Fortran-based [NOL97]. Functions [SSU97].

Gang [Jet97]. Garbage [ETY97]. General [WKC97]. Global [MR97, TNM^{+}97]. Grid [WM97].

Handling [Joh97]. Hardware [ABC^{+}97]. Hardware-Software [ABC^{+}97]. Hartree [KKCB97]. Hierarchy [HLK97]. High [ACM97, GBF^{+}97, GH97, HJG^{+}97, HJ97, Joh97, MT97, SdG^{+}97, TNM^{+}97, NOL97, RMCKB97]. High-end [GBF^{+}97]. High-fidelity [HJG^{+}97]. High-speed [Joh97]. Highly [BS97b]. HPF [LZ97]. Hydrodynamic [SS97]. Hydrodynamics [Fie97].

I/O [BDK97, KKCB97]. ILU [KK97]. Image [EGCL97]. Implementation [LZ97]. Implementations [BS97b, Fie97, WM97]. Implementing [WSP97]. Information [MFGD97]. Infrastructure [PBEK97]. Input [MR97]. Input/Output [MR97]. Instrumentation [Joh97]. Integrating [MT97]. Intel [GH97]. Interconnect [Cha97]. Interoperability [Lei97]. Issues [Lei97].

Jose [ACM97].

LAN [TNM^{+}97]. LAN/WAN [TNM^{+}97]. Languages [NSC97]. LANL [LLWB97]. Large [ETY97, SB97]. Large-scale [ETY97, SB97]. level [HJG^{+}97, KG97]. Libraries [ABE^{+}97]. Life [NOL97]. Limitations [CCvK97]. Linear [ABE^{+}97, LZ97, LSA97]. Load [HK97]. Long [TFS^{+}97]. Long-duration [TFS^{+}97]. Loop [VG97]. Loss [SSU97].

Machines [ETY97]. Management [KM97, LSA97, PBEK97]. Marangoni [CHMS97]. Marginal [SSU97]. Mark [ETY97]. Mark-sweep [ETY97]. Massively [CJK^{+}97, Fau97]. Matlab [MT97]. Matrix [KPS97]. Measuring [HLK97]. Memory [BS97a, HLK97, LLWB97, BW97, ETY97, Fie97, SB97]. Memory-centric [LLWB97]. Mergers [SS97]. Message [CLP97, Fie97]. Message-passing [CLP97, Fie97]. Messages [LMC97]. Metacomputing [WSP97]. Methods [BS97b, WM97]. Micro [GH97, HLK97]. Micro-Ops [GH97]. Mining [TOK97]. Model [NOL97, TFS^{+}97, WKC97, BW97]. Molecular [HK97, SB97]. MPI [LZ97]. MPMD [CCvK97]. MPP [CHMS97]. Multi [HJG^{+}97, LMC97, TNM^{+}97]. Multi-block [HJG^{+}97]. Multi-client [TNM^{+}97]. Multi-level [HJG^{+}97]. Multi-protocol [LMC97]. MultiMATLAB [MT97]. Multiprocessor [ABGS97]. Multiprocessors [HLK97]. Multiprogrammed [Jet97].

Nanotubes [SB97]. NASA [LEG^{+}97]. Network [WSP97]. Networked [MFGD97]. Networking [ACM97]. Neutron [SS97]. Ninf [TNM^{+}97]. Node [TOK97]. Noise [de 97]. Non [LSA97]. Non-linear [LSA97]. November [ACM97]. Numerical [LLC^{+}97].

O [BDK97, KKCB97]. Ocean [WKC97, TFS^{+}97]. Office [LEG^{+}97]. Online [Joh97]. Ops [GH97]. Optimization [Fie97, KKCB97, LSA97, WKC97]. ordering [VG97]. Origin2000 [LLWB97]. Overset [WM97]. Overview [ABE^{+}97].

Page [SSU97]. Parallel [ABE^{+}97, BDK97, BS97b, CLP97, CJK^{+}97, EGCL97, Fau97, HK97, KK97, KG97, KPS97, LSA97, LZ97, LEG^{+}97, MT97, NMM97, NSC97, TOK97, WKC97, WM97]. Parallelization [HJG^{+}97]. Parameters [BS97a]. PARDIS [KG97]. passing [CLP97, Fie97]. PASSION [KKCB97]. Pattern [MR97]. PC [TOK97]. PDE [PBEK97]. IEEE [ACM97]. Output [MR97]. WAN [TNM^{+}97]. Pentium [GH97]. Performance [ACM97, ABGS97, BS97a, CCvK97, GH97, HJ97, HLK97, Jet97, KM97, LLWB97, MT97, NSC97, NOL97, RMCKB97, SdG^{+}97, TNM^{+}97, BW97, WSP97]. Persistence [Lei97]. Phenomena [TFS^{+}97]. PLAPACK [ABE^{+}97]. Polymeric [Fau97]. Portable [BS97b, NSC97]. Porting [NOL97]. Pre [VG97]. Pre-fetching [VG97]. Preserving [HK97]. Pro [GH97]. Proceedings [ACM97]. Processing [MFGD97, TOK97]. Processors [GH97]. Programming [BW97]. Programs [BDK97, CLP97]. protocol [LMC97].

QCDSP [CJK^{+}97]. QoS [CC97]. Quality [LSA97]. Query [TOK97].

Radiation [SS97].

Radiation-hydrodynamic [SS97].

Rayleigh [CHMS97].

Rayleigh-Benard-Marangoni [CHMS97]. Re [VG97]. Re-ordering [VG97]. Real [CC97]. Real-time [CC97]. Recovery [NMM97]. Registration [EGCL97]. Remote [HJ97]. Replacement [SSU97]. Results [NOL97]. Run [VG97]. Run-time [VG97].

San [ACM97]. SC97 [ACM97, ACM97]. Scalable [ETY97, MFGD97, NMM97, SS97]. scale [CJK^{+}97, ETY97, SB97]. Scaling [Fie97]. Schedules [CC97]. Scheduling [Jet97]. Self [CC97]. Self-synchronizing [CC97]. Sensing [HJ97]. Service [WSP97]. SGI [LLWB97]. Shared [BS97a, ETY97, Fie97, SB97, BW97]. Shared-memory [ETY97, Fie97, SB97, BW97]. Simulation [BDK97, SB97]. Simulations [Fau97, HK97]. SMP [Cha97]. SMPs [LMC97, BW97]. SNIPE [MFGD97]. Software [ABC^{+}97, GBF^{+}97, GH97]. Solution [CHMS97]. Solutions [PBEK97]. Solver [LZ97]. Solvers [HJG^{+}97]. Sparse [KPS97]. speed [Joh97]. Spot [de 97]. Starfire [Cha97]. Stars [SS97]. Stencils [RMCKB97]. Strategy [NMM97]. Structure [SdG^{+}97]. Structures [Fau97]. Studies [SS97]. Study [BW97]. Studying [TFS^{+}97]. Supercomputer [CJK^{+}97]. Support [KM97, LSA97, TOK97]. Supporting [Lei97]. sweep [ETY97]. synchronizing [CC97]. System [GBF^{+}97, LSA97, TNM^{+}97, WSP97]. Systems [BDK97, Joh97, NMM97].

T3E [ABGS97, LLC^{+}97]. Tera [ABC^{+}97]. Teraflop [CJK^{+}97]. Teraflop-scale [CJK^{+}97]. TeraFLOPS [GH97]. Threshold [KK97]. Threshold-based [KK97]. time [CC97, VG97]. Tokamak [LLC^{+}97]. Tool [CLP97]. Topology [HK97]. Tuning [KM97]. Turbulence [LLC^{+}97].

USA [ACM97]. Using [CC97, HLK97, LZ97, SSU97, BW97].

Virtual [BS97a, Lei97].

Wavelet [EGCL97]. Wavelet-based [EGCL97]. Weather [WSP97].

Zeus [Fie97].

References

Alverson:1997:THS

[ABC^{+}97] Gail Alverson, Preston Briggs, Susan Coatney, Simon Kahan, and Rich Korry. Tera hardwaresoftware cooperation. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
ALVERSON/INDEX.HTM. ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Alpatov:1997:PPL

[ABE^{+}97] Philip Alpatov, Gregory Baker, H. Carter Edwards, John Gunnels, Greg Morrow, James Overfelt, and Robert van de Geijn. PLAPACK: Parallel linear algebra libraries design overview. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL http:

//www.supercomp.org/sc97/
proceedings/TECH/ALPATOV/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Anderson:1997:PAT

[ABGS97] Ed C. Anderson, Jeffrey P. Brooks, Charles M. Gassi, and Steven L. Scott. Performance analysis of the T3E multiprocessor. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL http:

//www.supercomp.org/sc97/
proceedings/TECH/ANDERSON/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

ACM:1997:SHP

[ACM97] ACM, editor. SC’97: High Performance Networking and Computing: Proceedings of the 1997 ACM/IEEE SC97 Conference: November 15–21, 1997, San Jose, California, USA. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-89791-985-8. LCCN ???? URL http:

//www.supercomp.org/sc97/
proceedings/. ACM SIGARCH

order number 415972. IEEE Computer Society Press order number RS00160.

Bagrodia:1997:PSP

[BDK97] Rajive Bagrodia, Stephen Docy, and Andy Kahn. Parallel simulation of parallel file systems and I/O programs. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
BAGRODIA/INDEX.HTM. ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Bilas:1997:ECP

[BS97a] Angelos Bilas and Jaswinder Pal Singh. The effects of communication parameters on end performance of shared virtual memory clusters. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
BILAS/INDEX.HTM. ACM SIGARCH

order number 415972. IEEE Computer Society Press order number RS00160.

Blackston:1997:HPE

[BS97b] David Blackston and Torsten Suel. Highly portable and efficient implementations of parallel adaptive N-body methods. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL http:

                                        //www.supercomp.org/sc97/
                                        proceedings/TECH/BLACKSTO/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Warren:1997:SPS

[BW97] Eugene D. Brooks III and Karen H. Warren. A study of performance on SMPs and distributed memory architectures using a shared-memory programming model. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

                                        URL http://www.supercomp.
                                        org/sc97/proceedings/TECH/
                                        BROOKS/INDEX.HTM.  ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Connelly:1997:FQR

[CC97] Kay H. Connelly and Andrew A. Chien. FM-QoS: Real-time communication using self-synchronizing schedules. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL http:

                                        //www.supercomp.org/sc97/
                                        proceedings/TECH/CONNELLY/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Chang:1997:EPL

[CCvK97] Chi-Chao Chang, Grzegorz Czajkowski, Thorsten von Eicken, and Carl Kesselman. Evaluating the performance limitations of MPMD communication. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
CHANG/INDEX.HTM. ACM SIGARCH

order number 415972. IEEE Computer Society Press order number RS00160.

Charlesworth:1997:SSI

[Cha97] Alan Charlesworth. The Starfire SMP interconnect. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
CHARLESW/INDEX.HTM. ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Carey:1997:MSR

[CHMS97] Graham F. Carey, Christopher Harle, Robert Mclay, and Spencer Swift. MPP solution of Rayleigh-Benard-Marangoni flow. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
CAREY/INDEX.HTM. ACM SIGARCH

order number 415972. IEEE Computer Society Press order number RS00160.

Christ:1997:QTS

[CJK^{+}97] Norman Christ, Chulwoo Jung, Adrian Kahler, Steven Kasow, Yubing Luo, Robert Mawhinney, Cheng-Zhong Sui, Pavlos Vranas, Alan Gara, John Parsons, Robert G. Edwards, Anthony D. Kennedy, Sten Hensen, Greg Kilcup, and Jim Sexton. QCDSP: A teraflop-scale massively parallel supercomputer. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

                                        URL http://www.supercomp.
                                        org/sc97/proceedings/TECH/
                                        CHRIST/INDEX.HTM.  ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Chen:1997:CCT

[CLP97] Yuqun Chen, Kai Li, and James S. Plank. CLIP: A checkpointing tool for messagepassing parallel programs. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

                                        URL http://www.supercomp.
                                        org/sc97/proceedings/TECH/
                                        CHEN/INDEX.HTM. ACM SIGARCH

order number 415972. IEEE Computer Society Press order number RS00160.

deLeeuw:1997:DCS

[de 97] Wim de Leeuw. Divide and conquer spot noise. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

                                        URL http://www.supercomp.
                                        org/sc97/proceedings/TECH/
                                        DELEEUW/INDEX.HTM.  ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

El-Ghazaw:1997:WBI

[EGCL97] A. El-Ghazaw, Prachya Chalermwa, and Jacqueline LeMoigne. Waveletbased image registration on parallel computers. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
ELGHAZAW/INDEX.HTM. ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Endo:1997:SMS

[ETY97] Toshio Endo, Kenjiro Taura, and Akinori Yonezawa. A scalable mark-sweep garbage collector on large-scale sharedmemory machines. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
ENDO/INDEX.HTM. ACM SIGARCH

order number 415972. IEEE Computer Society Press order number RS00160.

Faulon:1997:MPS

[Fau97] Jean-Loup M. Faulon. Massively parallel simulations of diffusion in dense polymeric structures. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
FAULON/INDEX.HTM.  ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Fiedler:1997:OSS

[Fie97] Robert A. Fiedler. Optimization and scaling of shared-memory and message-passing implementations of the Zeus hydrodynamics algorithm. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL http:

                                        //www.supercomp.org/sc97/
                                        proceedings/TECH/FIEDLER/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Greenberg:1997:SSA

[GBF^{+}97] David S. Greenberg, Ron Brightwell, Lee Ann Fisk, Arthur McCabe, and Rolf Riesen. A system software architecture for high-end computing. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL http:

                                        //www.supercomp.org/sc97/
                                        proceedings/TECH/GREENBER/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Greer:1997:HPS

[GH97] Bruce Greer and Greg Henry. High performance software on Intel Pentium Pro processors or micro-ops to TeraFLOPS. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

                                        URL http://www.supercomp.
                                        org/sc97/proceedings/TECH/
                                        GREER/INDEX.HTM. ACM SIGARCH

order number 415972. IEEE Computer Society Press order number RS00160.

Hawick:1997:DHP

[HJ97] Kenneth A. Hawick and Heath A. James. Distributed high performance computation for remote sensing. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
HAWICK/INDEX.HTM.  ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Hatay:1997:MLP

[HJG^{+}97] Ferhat F. Hatay, Dennis C. Jespersen, Guru P. Guruswamy, Yehia M. Rizk, Chansup Byun, and Ken Gee. A multi-level parallelization concept for highfidelity multi-block solvers. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
HATAY/INDEX.HTM. ACM SIGARCH

order number 415972. IEEE Computer Society Press order number RS00160.

Hegarty:1997:TPD

[HK97] David F. Hegarty and Tahar M.

Kechadi. Topology preserving dynamic load balancing for parallel molecular simulations. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL http:

//www.supercomp.org/sc97/
proceedings/TECH/HEGARTY/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Hristea:1997:MMH

[HLK97] Cristina Hristea, Daniel Lenoski, and John Keen. Measuring memory hierarchy performance of cache-coherent multiprocessors using micro benchmarks. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL http:

                                        //www.supercomp.org/sc97/
                                        proceedings/TECH/HRISTEA/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Jette:1997:PCG

[Jet97] Morris A. Jette. Performance characteristics of gang scheduling in multiprogrammed environments. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL http:

                                        //www.supercomp.org/sc97/
                                        proceedings/TECH/JETTE/INDEX.

HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Johnston:1997:HSD

[Joh97] William E. Johnston. Highspeed distributed data handling for online instrumentation systems. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL http:

                                        //www.supercomp.org/sc97/
                                        proceedings/TECH/JOHNSTON/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Keahey:1997:PCB

[KG97] Katarzyna Keahey and Dennis Gannon. PARDIS: CORBAbased architecture for applicationlevel parallel distributed computation. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL http:

//www.supercomp.org/sc97/
proceedings/TECH/KEAHEY/INDEX.

HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Karypis:1997:PTB

[KK97] George Karypis and Vipin Kumar. Parallel threshold-based ILU factorization. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
KARYPIS/INDEX.HTM.  ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Kandaswamy:1997:OEH

[KKCB97] Meenakshi A. Kandaswamy, Mahmut T. Kandemir, Alok N. Choudhary, and David E. Bernholdt. Optimization and evaluation of Hartree–Fock applications’ I/O with PAS-SION. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL http:

//www.supercomp.org/sc97/
proceedings/TECH/KANDASWA/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Karavanic:1997:EMS

[KM97] Karen L. Karavanic and Barton P. Miller. Experiment management support for performance tuning. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
KARAVANI/INDEX.HTM. ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Kotlyar:1997:CPC

[KPS97] Vladimir Kotlyar, Keshav Pingali, and Paul Vinson Stodghill. Compiling parallel code for sparse matrix applications. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
KOTLYAR/INDEX.HTM.  ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Lyster:1997:PCN

[LEG^{+}97] M P. Lyster, K. Ekers, J. Guo, M. Harber, D. Lamich, J. W. Larson, R. Lucchesi, R. Rood, S. Schubert, W. Sawyer, M. Sienkiewicz, A. da Silva, J. Stobie, L. L. Takacs, R. Todling, J. Zero, C. H. Q. Ding, and R. Ferraro. Parallel computing at NASA Data Assimilation Office. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
LYSTER/INDEX.HTM.  ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Leigh:1997:IDF

[Lei97] Jason Leigh. Issues in the design of a flexible distributed architecture for supporting persistence and interoperability in collaborative virtual environments. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
LEIGH/INDEX.HTM. ACM SIGARCH

order number 415972. IEEE Computer Society Press order number RS00160.

Lynch:1997:NTT

[LLC^{+}97] Vickie E. Lynch, Jean-Noel Leboeuf, Benjamin A. Carreras, Juan Diego Alvarez, and Luis Garcia. Numerical Tokamak turbulence calculations on the CRAY T3E. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
LYNCH/INDEX.HTM. ACM SIGARCH

order number 415972. IEEE Computer Society Press order number RS00160.

Lubeck:1997:PES

[LLWB97] Olaf M. Lubeck, Yong Luo, Harvey J. Wasserman, and Federico Bassetti. Performance evaluation of the SGI Origin2000: A memory-centric characterization of LANL ASCI applications. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
LUBECK/INDEX.HTM.  ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Lumetta:1997:MPA

[LMC97] Steven S. Lumetta, Alan Mainwaring, and David E. Culler. Multi-protocol active messages on a cluster of SMPs. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL http:

                                        //www.supercomp.org/sc97/
                                        proceedings/TECH/LUMETTA/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Lewis:1997:PNL

[LSA97] Andrew Lewis, Rod Simpson, and David Abramson. Parallel non-linear optimization: Toward the design of a decision support system for air quality management. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

                                        URL http://www.supercomp.
                                        org/sc97/proceedings/TECH/
                                        LEWIS/INDEX.HTM. ACM SIGARCH

order number 415972. IEEE Computer Society Press order number RS00160.

Li:1997:EHC

[LZ97] Konming Gary Li and Nabil M. Zamel. An evaluation of HPF compilers and the implementation of a parallel linear equation solver using HPF and MPI. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

                                        URL http://www.supercomp.
                                        org/sc97/proceedings/TECH/

LI/INDEX.HTM. ACM SIGARCH

order number 415972. IEEE Computer Society Press order number RS00160.

Moore:1997:SNI

[MFGD97] Keith Moore, Graham E. Fagg, Al Geist, and Jack Dongarra. Scalable networked information processing environment (SNIPE. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
MOORE/INDEX.HTM. ACM SIGARCH

order number 415972. IEEE Computer Society Press order number RS00160.

Madhyastha:1997:EGI

[MR97] Tara M. Madhyastha and Daniel A. Reed. Exploiting global input/output access pattern classification. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
MADHYAST/INDEX.HTM. ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Menon:1997:MIM

[MT97] Vijay Menon and Anne E. Trefethen. MultiMATLAB: Integrating Matlab with high performance parallel computing. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
MENON/INDEX.HTM. ACM SIGARCH

order number 415972. IEEE Computer Society Press order number RS00160.

Naik:1997:CSS

[NMM97] Vijay K. Naik, Samuel P. Midkiff, and Jose E. Moreira. A checkpointing strategy for scalable recovery on distributed parallel systems. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

                                        URL http://www.supercomp.
                                        org/sc97/proceedings/TECH/
                                        NAIK/INDEX.HTM. ACM SIGARCH

order number 415972. IEEE Computer Society Press order number RS00160.

Nucciarone:1997:NLD

[NOL97] Jeffrey J. Nucciarone, Yusuf Ozyoruk, and Lyle N. Long. New life in dusty decks: Results of porting a CM Fortranbased aeroacoustic model to High Performance Fortran. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

                                        URL http://www.supercomp.
                                        org/sc97/proceedings/TECH/
                                        NUCCIARO/INDEX.HTM. ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Ngo:1997:PPD

[NSC97] Ton A. Ngo, Lawrence Snyder, and Bradford Chamberlain. Portable performance of data parallel languages. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL

                                        http://www.supercomp.org/
                                        sc97/proceedings/TECH/NGO/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Parashar:1997:CDM

[PBEK97] M. Parashar, J. C. Browne, C. Edwards, and K. Klimkowski. A common data management infrastructure for adaptive algorithms for PDE solutions. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL http:

//www.supercomp.org/sc97/
proceedings/TECH/PARASHAR/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Roth:1997:CSH

[RMCKB97] Gerald Roth, John Mellor-Crummey, Ken Kennedy, and R. Gregg Brickner. Compiling stencils in High Performance Fortran. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
ROTH/INDEX.HTM. ACM SIGARCH

order number 415972. IEEE Computer Society Press order number RS00160.

Srivastava:1997:MDS

[SB97] Deepak Srivastava and Stephen T.

Barnard. Molecular dynamics simulation of large-scale carbon nanotubes on a shared-memory architecture. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL http:

//www.supercomp.org/sc97/

                                        proceedings/TECH/SRIVASTA/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Stout:1997:ABH

[SdG^{+}97] Q. F. Stout, D. deZeeuw, T. Gombossi, C. Groth, H. Marshall, and K. Powell. Adaptive blocks: A high performance data structure. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

                                        URL http://www.supercomp.
                                        org/sc97/proceedings/TECH/
                                        STOUT/INDEX.HTM. ACM SIGARCH

order number 415972. IEEE Computer Society Press order number RS00160.

Saylor:1997:SHR

[SS97] Paul E. Saylor and Frank D. Swesty. Scalable, hydrodynamic, and radiation-hydrodynamic studies of neutron stars mergers. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

                                        URL http://www.supercomp.
                                        org/sc97/proceedings/TECH/
                                        SAYLOR/INDEX.HTM.  ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Sharma:1997:PRU

[SSU97] Shamik Das Sharma, Joel Saltz, and Manuel Ujaldon. Page replacement using marginal loss functions. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

                                        URL http://www.supercomp.
                                        org/sc97/proceedings/TECH/

SHARMA/INDEX.HTM.  ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Tobis:1997:FAO

[TFS^{+}97] Michael Tobis, Ian T. Foster, Chad M. Schafer, Robert M. Jacob, and John R. Anderson. FOAM: An atmosphereocean climate model for studying long-duration phenomena. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
TOBIS/INDEX.HTM. ACM SIGARCH

order number 415972. IEEE Computer Society Press order number RS00160.

Takefusa:1997:MCL

[TNM^{+}97] Atsuko Takefusa, Umpei Nagashima, Satoshi Matsuoka, Hirotaka Ogawa, Hidemoto Nakada, Satoshi Sekiguchi, Hiromitsu Takagi, and Mitsuhisa Sato. Multi-client LAN/WAN performance analysis of Ninf: A high performance global computing system. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL http:

//www.supercomp.org/sc97/
proceedings/TECH/TAKEFUSA/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Tamura:1997:PDP

[TOK97] Takayuki Tamura, Masato Oguchi, and Masaru Kitsuregawa. Parallel database processing on a 100 node PC cluster: Cases for decision support query processing and data mining. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

                                        URL http://www.supercomp.
                                        org/sc97/proceedings/TECH/
                                        TAMURA/INDEX.HTM.  ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Vajracharya:1997:LRO

[VG97] Suvas Vajracharya and Dirk Grunwald. Loop re-ordering and pre-fetching during runtime. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL http:

                                        //www.supercomp.org/sc97/
                                        proceedings/TECH/VAJRACHA/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Wang:1997:OPO

[WKC97] Ping Wang, Daniel S. Katz, and Yi Chao. Optimization of a parallel ocean general circulation model. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

                                        URL http://www.supercomp.
                                        org/sc97/proceedings/TECH/
                                        WANG/INDEX.HTM. ACM SIGARCH

order number 415972. IEEE Computer Society Press order number RS00160.

Wissink:1997:PID

[WM97] Andrew M. Wissink and Robert L. Meakin. On parallel implementations of dynamic overset grid methods. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL http:

//www.supercomp.org/sc97/
proceedings/TECH/WISSINK/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Wolski:1997:IPF

[WSP97] Rich Wolski, Neil Spring, and Chris Peterson. Implementing a performance forecasting system for metacomputing: The network weather service. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ????

URL http://www.supercomp.
org/sc97/proceedings/TECH/
WOLSKI/INDEX.HTM.  ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.